IN what time will interest on a given sum of at 5 1/3 amount to one third of the sum?
Answer by ankor@dixie-net.com(6567)

IN what time will interest on a given sum of at 5 1/3 amount to one third of the sum?
:
Using p*r*t = i
Where
p = 3
r = .0533
t = time
i = 1
:
3*.0533*t = 1
.0533t =
.0533t = .333
t =
t = 6.247 yrs
We have to say 7 yrs if they pay annually
:
:
Check on calc: enter: 3*.0533*6.247 = .999 ~ 1